I have seen this problem covered in a lot of other posts, but couldn't find the solution to my particular problem.
Started getting intermittent switch failure on Balrog target, then it finally stopped responding altogether. During switch test I found the switch would activate intermittently when I would hold the Balrog target 2mm from closed (blocking the ring ramp). When fully open (not in use) the switch would once again work intermittently. It was only when the Balrog target was fully closed it would not respond.
I replaced the switch and diode, took off the shrink wrap to inspect the cable for any breaks. I didn't see any so I re-shrink wrapped it. Tested the switch in test mode before attaching it back to Balrog and it worked great. I twisted and moved the cabling during the test and still got consistent switch registrations. Put it all back together got it to respond for one round of FoTR, put the glass back on and it is doing the same thing as before. The switch only responds when Balrog is 2mm from fully closed.
Before I replace the wiring harness altogether I was wondering if anyone has any suggestions for anything else I should look at first?